# MILo: Mesh-In-the-Loop Gaussian Splatting for Detailed and Efficient Surface Reconstruction

### v1

@@ Abstract.
A presentation video of our approach is available at https://youtu.be/_SGNhhNz0fE .
@@
submissionid: 1884journal: TOG
Figure 1. Mesh-in-the-Loop Gaussian Splatting. Our method introduces a novel differentiable mesh extraction framework that operates during the optimization of 3D Gaussian Splatting representations. At every training iteration, we differentiably extract a mesh—including both vertex locations and connectivity—only from Gaussian parameters. This enables gradient flow from the mesh to Gaussians, allowing us to promote bidirectional consistency between volumetric (Gaussians) and surface (extracted mesh) representations. This approach guides Gaussians toward configurations better suited for surface reconstruction, resulting in higher quality meshes with significantly fewer vertices. In this example, our method reconstructs the entire bicycle scene—including background—with almost 10 times fewer vertices than previous methods while preserving fine geometric details. Our framework can be plugged into any Gaussian splatting representation, increasing performance while generating an order of magnitude fewer mesh vertices. MILo makes the reconstructions more practical for downstream applications like physics simulations and animation.
@@ 1. Introduction
This postprocessing step collapses complex volumetric information into a surface, often leading to the loss of fine or semitransparent structures, and can introduce artifacts in the extracted mesh. Moreover, Gaussian Splatting and NeRF-based methods are known to adjust their opacity and view-dependent colors independently of the geometry. This allows them to fit the training images more precisely, but often at the expense of geometric consistency. “Cheating” leads to hallucinated structures such as floaters or cavities, which are particularly hard to resolve during mesh extraction. Fixing them is non-trivial, as the underlying volumetric representation has already absorbed these inconsistencies in its optimization. Several works have highlighted this issue in challenging settings—such as low-information (Warburg et al., 2023; Yang et al., [n. d.]; Gomez et al., 2025; Guédon et al., 2025) or high specularities (Verbin et al., 2022)—yet a general and robust solution remains elusive.
@@
Moreover, many existing methods (Yariv et al., 2021; Peng Wang et al., 2021; Huang et al., 2024; Zhang et al., 2024) focus on object-centric reconstructions, evaluating only the quality of the foreground geometry. Similarily, benchmarks like DTU (Jensen et al., 2014) or Tanks and Temples (Knapitsch et al., 2017) offer ground truth solely for foreground objects, discouraging full-scene reconstructions that have backgrounds.
@@
This setup enables direct backpropagation from the mesh to the Gaussian parameters, allowing us to impose consistency between the surface mesh, volumetric field, and input data. Thanks to the Pivots, Gaussians serve as an implicit parameterization of an explicit mesh, progressively optimized to yield high-quality geometry, vertex positions, and connectivity.
@@
We achieve state-of-the-art geometric quality while using an order of magnitude fewer vertices than competing methods, making our surfaces far more practical for downstream applications. To evaluate mesh quality in background regions where no ground-truth geometry exists, we propose a new proxy metric based on mesh rendering: training a neural color field over the mesh and comparing rendered images across methods. This decoupling rendering fidelity from mesh resolution and helps us to evaluate reconstruction quality across the entire scene.
Our contributions are:
@@
  • We design an evaluation protocol, Mesh-Based Novel View synthesis, to assess full-scene geometry, even in the absence of ground-truth 3D models.
We provide detailed definitions of 3D Gaussian Splatting and Delaunay Triangulation in the Supplemental.
@@ 2. Related Work
Recently, 3D Gaussian Splatting (Kerbl et al., 2023) has redefined the NVS landscape by enabling faster, high-quality synthesis from point-based primitives. This sparked a wave of follow-up work aiming to improve efficiency and scalability, including Mip-Splatting (Yu et al., 2024), Mini-Splatting v1 and v2 (Fang and Wang, 2024a, b), MCMC-Gaussians (Kheradmand et al., 2024), BetaSplatting (Liu et al., 2025), and Taming 3DGS (Mallick et al., 2024). Additional work like 3DGUT (Wu et al., 2024) and 3DGRT (Moenne-Loccoz et al., 2024) extend these models to arbitrary camera systems and differentiable ray tracing, respectively.
@@
Recovering explicit surface meshes from images only is a longstanding challenge. Recent volumetric approaches relying on implicit functions (Peng Wang et al., 2021; Yariv et al., 2021; Li et al., 2023) proved efficient for reconstructing accurate surfaces, but generally require very long training times (often exceeding 24 hours) before mesh extraction can be performed. Moreover, these volumetric representations typically cannot achieve real-time rendering, limiting their practical applications. Some approaches address this by first optimizing the volumetric representation, extracting a mesh, and then baking the rendering capabilities into the mesh to enable real-time rendering (Yariv et al., 2023; Reiser et al., 2024), but these still suffer from the lengthy optimization times required for the initial mesh extraction. Other methods based on discrete structures (Munkberg et al., 2022; Shen et al., 2023) also proved efficient to some extent, but focus on individual objects and are not scalable to large complex scenes.
@@
In contrast, recent methods relying on Gaussian Splatting offer promising alternatives: 2DGS (Huang et al., 2024), RaDe-GS (Zhang et al., 2024), NeuSG (Chen et al., 2023), VCR-GauS (Chen et al., 2024), and Quadratic Gaussian Splatting (Zhang et al., 2024) explore different rendering or regularization strategies that rely on Truncated Signed Distance Fields (TSDF), while SuGaR (Guédon and Lepetit, 2024b) and Gaussian Opacity Fields (Yu et al., 2024) propose their own mesh extraction procedures.
@@
However, all existing methods perform mesh extraction after optimization, treating it as a post-processing step, or adopt a two-stage approach where they first optimize a volumetric representation, extract a mesh, and then refine it with differentiable rendering (Yariv et al., 2023; Guédon and Lepetit, 2024b, a). The latter approach still suffers from the same initial extraction issues, and additionally constrains the mesh topology during refinement, which stays fixed after extraction. This separation introduces potential inconsistencies: there is no guarantee that the extracted surface accurately reflects the underlying volumetric representation. Since Gaussians are continuous by nature, naive isosurfacing often results in geometric artifacts such as over-inflation or erosion. These artifacts are particularly noticeable around thin structures (see Fig. 1). Moreover, these methods typically produce overly dense meshes (up to tens of millions of vertices) that are difficult to scale to large scenes. In contrast, our approach integrates mesh extraction directly into the optimization loop, enabling consistent supervision of surface geometry and yielding compact, high-fidelity scalable meshes.
Voronoi &\&& Delaunay-based methods.
@@ 3. Overview
Our method aims to address the fundamental challenge of extracting high-quality surface meshes from 3D Gaussian Splatting representations. Rather than treating mesh extraction as a post-processing step, we integrate it directly into the optimization process and ensure that both the volumetric and surface representations are consistent. By extracting a mesh at every training iteration and back-propagating gradients to the Gaussians, we guide them toward configurations optimized for accurate surface extraction. See Fig. 2 for a visual overview our method.
@@ 4.1. Sampling Delaunay Vertices from Gaussians
Table 2. Quantitative comparison on the Tanks & Temples Dataset (Knapitsch et al., 2017). We report the F1-score and average optimization time. All results are evaluated with the official evaluation scripts. $\text{Ours}_{\text{X}}$ uses the differentiable rasterizer from X. VCR-GauS (Chen et al., 2024) relies on a pre-trained normal estimation model. Best results for implicit methods are highlighted in blue; best results for explicit methods are highlighted in red. Our method achieves the best F1 score among explicit representations, even though it outputs lighter meshes than other approaches.
@@
NeuS Geo-Neus Neuralangelo SuGaR 3DGS 2DGS GOF RaDe-GS QGS VCR-GauS $\text{Ours}_{\text{RaDe-GS}}$ (base) $\text{Ours}_{\text{GOF}}$ (base) $\text{Ours}_{\text{RaDe-GS}}$ (dense)
Barn 0.29 0.33 0.70 0.14 0.13 0.36 0.51 0.43 0.43 0.62 0.55 0.59 0.64
Caterpillar 0.29 0.26 0.36 0.16 0.08 0.23 0.41 0.32 0.31 0.26 0.38 0.39 0.38
Courthouse 0.17 0.12 0.28 0.08 0.09 0.13 0.28 0.21 0.26 0.19 0.28 0.29 0.31
Ignatius 0.83 0.72 0.89 0.33 0.04 0.44 0.68 0.69 0.79 0.61 0.73 0.78 0.76
Meetingroom 0.24 0.20 0.32 0.15 0.01 0.16 0.28 0.25 0.25 0.19 0.25 0.26 0.28
Truck 0.45 0.45 0.48 0.26 0.19 0.26 0.59 0.51 0.60 0.52 0.60 0.62 0.59
Mean 0.38 0.35 0.50 0.19 0.09 0.30 0.46 0.40 0.44 0.40 0.47 0.49 0.49
Time ¿24h ¿24h ¿24h 73 m 7.9 m 12.3  m 69m 11.5 m 120 m 53 m 50 m 150 m 110 m
@@
  • Dense model: We also sample a set of Gaussians using Importance-weighted sampling (Fang and Wang, 2024b), but we do not remove the other Gaussians from the scene: We maintain the large set of Gaussians (typically between 2M and 5M) for Gaussian Splatting but use only the sampled ones as pivots for generating the Delaunay vertices. The number of Delaunay vertices is approximately the same as the base model, still resulting in a lightweight mesh; however, having more Gaussians helps in learning better SDF values at these vertices. This model results in longer optimization times, but better performances.
@@ 4.2. Computing Signed Distance Values
We propose a simple yet effective strategy: augmenting each Gaussian $\mathcal{G}_{k}$ with 9 optimizable SDF values $f_{k}\in\mathbb{R}^{9}$—one for each Delaunay vertex. Importantly, these SDF values are decoupled from the Gaussians’ other parameters—such as opacity, scale, and rotation—allowing for localized control over the extracted isosurface level. This decoupling proves highly beneficial for accurately capturing fine surface details and ensuring strong consistency between the extracted mesh and the underlying volumetric representation. For an even faster convergence, we devise a custom initialization algorithm: please refer to the supplemental for more details. An illustration of our approach is provided in Fig 3.
@@ 5.3. Regularization
To address this issue, we propose a regularization term ${\mathcal{}L}_{\text{erosion}}$ that prevents erosion and ensures that all geometry is preserved:
@@
where $G_{\text{Del}}$ is the set of Gaussians sampled for the Delaunay triangulation and $f_{\mu_{g}}$ is the SDF value of the center of the $g$-th Gaussian. As mentioned in Section 4, the centers of all Gaussians sampled for the triangulation are used as Delaunay vertices: this regularization term aims to include the centers of these Gaussians inside the surface by encouraging their SDF values to become negative, preventing erosion in all areas where at least one Gaussian has been sampled.
@@
To prevent such artifacts inside the mesh, we introduce a novel loss based on a feedback loop between the SDF values and the extracted mesh. This loss aims to enforce occluded points located inside the mesh to have negative SDF values. Specifically, after extracting the surface using the Gaussians and their SDF values as described in section 4.3,
we use the mesh to build an occupancy label $o_{p}\in\{0,1\}$ for each Delaunay site $p$, indicating whether these points are inside or outside the visible portion of the surface mesh as described below. We finally enforce the Delaunay vertices labeled as inside to have negative SDF values, without enforcing any additional constraint for Delaunay vertices labeled as outside—as these points already receive supervision from the renderings.
@@
Figure 5. Qualitative comparison of surface reconstruction results on the Tanks and Temples dataset. Our method produces meshes of full scenes that better fit the target surfaces, with reduced artifacts such as cavities (three left images). Note how our approach effectively addresses the erosion problem that plagues previous methods (two right images), while producing lighter meshes.
@@
implicit NeRF (Mildenhall et al., 2020) 1.90 1.60 1.85 0.58 2.28 1.27 1.47 1.67 2.05 1.07 0.88 2.53 1.06 1.15 0.96 1.49 $>\text{12h}$
VolSDF (Yariv et al., 2021) 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86 $>\text{12h}$
NeuS (Peng Wang et al., 2021) 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84 $>\text{12h}$
Neuralangelo (Li et al., 2023) 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61 $>\text{12h}$
explicit 3D GS (Kerbl et al., 2023) 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96 5.2m

### v2

@@ Abstract.
The code for our approach and an online gallery are available at https://anttwo.github.io/milo/ .
@@
submissionid: 1884journal: TOGjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 6article: 1publicationmonth: 12doi: 10.1145/3763339ccs: Computing methodologies Reconstructionccs: Computing methodologies Mesh modelsccs: Computing methodologies Point-based modelsccs: Computing methodologies Shape representationsccs: Computing methodologies Renderingccs: Computing methodologies Image-based rendering
Figure 1. Mesh-in-the-Loop Gaussian Splatting. Our method introduces a novel differentiable mesh extraction framework that operates during the optimization of 3D Gaussian Splatting representations. At every training iteration, we differentiably extract a mesh—including both vertex locations and connectivity—directly from Gaussian parameters. This enables gradient flow from the mesh to Gaussians, allowing us to promote bidirectional consistency between volumetric (Gaussians) and surface (extracted mesh) representations. This approach guides Gaussians toward configurations better suited for surface reconstruction, resulting in higher quality meshes with significantly fewer vertices. In this example, our method reconstructs the entire bicycle scene—including background—with almost 10 times fewer vertices than previous methods while preserving fine geometric details. Our framework can be plugged into any Gaussian splatting representation, increasing performance while generating an order of magnitude fewer mesh vertices. MILo makes the reconstructions more practical for downstream applications like physics simulations and animation.
@@ 1. Introduction
This postprocessing step collapses complex volumetric information into a surface, often leading to the loss of fine or semitransparent structures, and can introduce artifacts in the extracted mesh. Moreover, Gaussian Splatting and NeRF-based methods are known to adjust their opacity and view-dependent colors independently of the geometry. This allows them to fit the training images more precisely, but often at the expense of geometric consistency. “Cheating” leads to hallucinated structures such as floaters or cavities, which are particularly hard to resolve during mesh extraction. Fixing them is non-trivial, as the underlying volumetric representation has already absorbed these inconsistencies in its optimization. Several works have highlighted this issue in challenging settings—such as low-information (Warburg et al., 2023; Yang et al., 2023; Gomez et al., 2025; Guédon et al., 2025) or high specularities (Verbin et al., 2022)—yet a general and robust solution remains elusive.
@@
Moreover, many existing methods (Yariv et al., 2021; Wang et al., 2021; Huang et al., 2024; Zhang et al., 2024a) focus on object-centric reconstructions, evaluating only the quality of the foreground geometry. Similarly, benchmarks like DTU (Jensen et al., 2014) or Tanks and Temples (Knapitsch et al., 2017) offer ground truth solely for foreground objects, discouraging full-scene reconstructions that have backgrounds.
@@
This setup enables direct gradient backpropagation from the mesh to the Gaussian parameters, allowing us to impose consistency between the surface mesh, volumetric field, and input data. Thanks to the Pivots, Gaussians serve as an implicit parameterization of an explicit mesh, progressively optimized to yield high-quality geometry, vertex positions, and connectivity.
@@
We achieve state-of-the-art geometric quality while using an order of magnitude fewer vertices than competing methods, making our surfaces far more practical for downstream applications. Inspired in previous works which focus on surface-based view synthesis methods (Reiser et al., 2024) we use a proxy metric based on mesh rendering to evaluate mesh quality in background regions where no ground-truth geometry exists, training a neural color field over the mesh and comparing rendered images across methods. This decoupling rendering fidelity from mesh resolution and helps us to evaluate reconstruction quality across the entire scene. Our work makes the following contributions.
@@
  • We adapt a metric introduced in the context of surface-based view synthesis and use it as an evaluation protocol to assess full-scene geometry, even in the absence of ground-truth 3D models. We use this protocol to demonstrate that our method achieves state-of-the-art results in mesh quality and compactness across multiple complex 3D scenes,
We provide detailed definitions of 3D Gaussian Splatting and Delaunay Triangulation in the supplementary material.
@@ 2. Related Work
Recently, 3D Gaussian Splatting (Kerbl et al., 2023) has redefined the NVS landscape by enabling faster, high-quality synthesis from point-based primitives. This sparked a wave of follow-up work aiming to improve efficiency and scalability, including Mip-Splatting (Yu et al., 2024a), Mini-Splatting v1 and v2 (Fang and Wang, 2024a, b), MCMC-Gaussians (Kheradmand et al., 2024), BetaSplatting (Liu et al., 2025), and Taming 3DGS (Mallick et al., 2024). Additional work like 3DGUT (Wu et al., 2024) and 3DGRT (Moenne-Loccoz et al., 2024) extend these models to arbitrary camera systems and differentiable ray tracing, respectively.
@@
Recovering explicit surface meshes exclusively from images is a longstanding challenge. Well-established volumetric approaches relying on implicit functions (Wang et al., 2021; Yariv et al., 2021; Li et al., 2023) proved efficient for reconstructing accurate surfaces. Training an implicit function is a strong regularization that allows to obtain smooth surfaces, thus these methods excel at extracting the foreground of object-centric scenes.
However, these methods generally require very long training times (often exceeding 24 hours) before mesh extraction can be performed. The success of Gaussian Splatting has naturally inspired an emerging line of work that trains splats jointly with implicit functions, notably Gaussian-UDF (Li et al., 2025), GS-Pull (Zhang et al., 2024b), and GSDF (Yu et al., 2024b). By combining these two representations, such methods mitigate the time bottleneck by querying the implicit function through strategies that exploit the positions of the Gaussian splats. While this improves training efficiency, it also exacerbates scalability issues, since two representations must be optimized simultaneously. Moreover, although neural SDFs excel at reconstructing isolated objects, their expressivity remains bounded by the MLP architecture (Li et al., 2025; Zhang et al., 2024b; Yu et al., 2024b). Consequently, while these approaches can in principle be applied to more complex scenes, they tend to underperform beyond single-object or foreground-dominant settings.
Moreover, these volumetric representations typically cannot achieve real-time rendering, limiting their practical applications. Some approaches address this by first optimizing the volumetric representation, extracting a mesh, and then baking the rendering capabilities into the mesh to enable real-time rendering (Yariv et al., 2023; Reiser et al., 2024). Thus, these methods are split into two stages, which prevents any refinement of the mesh, while still exhibiting lengthy optimization times.
Other methods based on discrete structures (Munkberg et al., 2022; Shen et al., 2023) also proved efficient to some extent, but focus on individual objects and are not scalable to large complex scenes.
@@
Recent methods relying on Gaussian Splatting offer promising efficient alternatives (Wang et al., 2023; Dai et al., 2024) namely: 2DGS (Huang et al., 2024), RaDe-GS (Zhang et al., 2024), NeuSG (Chen et al., 2023), VCR-GauS (Chen et al., 2024), and Quadratic Gaussian Splatting (Zhang et al., 2024a) explore different rendering or regularization strategies that rely on Truncated Signed Distance Fields (TSDF), while SuGaR (Guédon and Lepetit, 2024b) and Gaussian Opacity Fields (Yu et al., 2024) propose their own mesh extraction procedures.
@@
All previously highlighted methods perform mesh extraction after optimization, treating it as a post-processing step. Some of these go slightly further, adopting a two-stage approach where they first optimize a volumetric representation, extract a mesh, and then refine it with differentiable rendering (Yariv et al., 2023; Guédon and Lepetit, 2024b, a). This line of work, however, neither avoids the initial extraction issues nor adjusts the mesh topology, which stays fixed during the refinement process. This separation, in both cases, introduces potential inconsistencies: there is no guarantee that the extracted surface accurately reflects the underlying volumetric representation. Since Gaussians are continuous by nature, naive isosurfacing often results in geometric artifacts such as over-inflation or erosion. These artifacts are particularly noticeable around thin structures (see Fig. 1). Moreover, all of these methods typically produce overly dense meshes (up to tens of millions of vertices) that are difficult to scale to large scenes.
In contrast, our method efficiently leverages Gaussians to scale to full-scene reconstruction, and is the first to integrate mesh extraction directly into the optimization loop. This design naturally enables the refinement of both vertex positions and mesh topology, ensuring consistency with the underlying volumetric representation (Gaussian splats).
Voronoi &\& Delaunay-based methods.
@@ 3. Overview
Our method aims to address the fundamental challenge of extracting high-quality surface meshes from 3D Gaussian Splatting representations. Rather than treating mesh extraction as a post-processing step, we integrate it directly into the optimization process and ensure that both the volumetric and surface representations are consistent. By extracting a mesh at every training iteration and back-propagating gradients to the Gaussians, we guide them toward configurations optimized for accurate surface extraction. See Fig. 2 for a visual overview of our method.
@@ 4.1. Sampling Delaunay Vertices from Gaussians
Table 2. Quantitative comparison on the Tanks & Temples Dataset (Knapitsch et al., 2017). We report the F1-score and average optimization time. All results are evaluated with the official evaluation scripts. $\text{Ours}_{\text{X}}$ uses the differentiable rasterizer from X. VCR-GauS (Chen et al., 2024) relies on a pre-trained normal estimation model. Best results for implicit methods are highlighted in blue; best results for explicit methods are highlighted in red. When a method fails (OOM), we report the mean of the successful scenes, denoted by a superscript asterisk ($x^{*}$). Our method achieves the best F1 score among explicit representations, even though it outputs lighter meshes than other approaches.
@@
Gaussian UDF GS-Pull GSDF NeuS Geo-Neus Neuralangelo SuGaR 3DGS 2DGS GOF RaDe-GS QGS VCR-GauS $\text{Ours}_{\text{RaDe-GS}}$ (base) $\text{Ours}_{\text{GOF}}$ (base) $\text{Ours}_{\text{RaDe-GS}}$ (dense)
Barn 0.27 0.60 0.16 0.29 0.33 0.70 0.14 0.13 0.36 0.51 0.43 0.43 0.62 0.55 0.59 0.64
Caterpillar 0.17 0.37 0.11 0.29 0.26 0.36 0.16 0.08 0.23 0.41 0.32 0.31 0.26 0.38 0.39 0.38
Courthouse 0.03 0.16 OOM 0.17 0.12 0.28 0.08 0.09 0.13 0.28 0.21 0.26 0.19 0.28 0.29 0.31
Ignatius 0.38 0.71 0.34 0.83 0.72 0.89 0.33 0.04 0.44 0.68 0.69 0.79 0.61 0.73 0.78 0.76
Meetingroom 0.17 0.22 0.01 0.24 0.20 0.32 0.15 0.01 0.16 0.28 0.25 0.25 0.19 0.25 0.26 0.28
Truck 0.30 0.52 0.25 0.45 0.45 0.48 0.26 0.19 0.26 0.59 0.51 0.60 0.52 0.60 0.62 0.59
Mean 0.22 0.43 $0.18^{*}$ 0.38 0.35 0.50 0.19 0.09 0.30 0.46 0.40 0.44 0.40 0.47 0.49 0.49
Time 90m 37.6m 70m ¿24h ¿24h ¿24h 73 m 7.9 m 12.3  m 69m 11.5 m 120 m 53 m 50 m 150 m 110 m
@@
  • Dense model: We also sample a set of Gaussians using Importance-weighted sampling (Fang and Wang, 2024b), but we do not remove the other Gaussians from the scene: We maintain the large set of Gaussians (typically between 2M and 5M) for Gaussian Splatting but use only the sampled ones as pivots for generating the Delaunay vertices. The number of Delaunay vertices is approximately the same as the base model, still resulting in a lightweight mesh; however, having more Gaussians helps in learning better SDF values at these vertices. This model results in longer optimization times, but better performance.
@@ 4.2. Computing Signed Distance Values
We propose a simple yet effective strategy: augmenting each Gaussian $\mathcal{G}_{k}$ with 9 optimizable SDF values $f_{k}\in\mathbb{R}^{9}$—one for each Delaunay vertex. Importantly, these SDF values are decoupled from the Gaussians’ other parameters—such as opacity, scale, and rotation—allowing for localized control over the extracted isosurface level. This decoupling proves highly beneficial for accurately capturing fine surface details and ensuring strong consistency between the extracted mesh and the underlying volumetric representation. For an even faster convergence, we devise a custom initialization algorithm: please refer to the supplementary material for more details. An illustration of our approach is provided in Fig 3.
@@ 5.3. Regularization
To address this issue, we propose a regularization term ${\mathcal{}L}_{\text{erosion}}$ that encourages geometry to be preserved:
@@
where $G_{\text{Del}}$ is the set of Gaussians sampled for the Delaunay triangulation and $f_{\mu_{g}}$ is the SDF value of the center of the $g$-th Gaussian. As mentioned in Section 4, the centers of all Gaussians sampled for the triangulation are used as Delaunay vertices: this regularization term aims to include the centers of these Gaussians inside the surface by encouraging their SDF values to become negative, preventing erosion in all areas where at least one Gaussian has been sampled. ${\mathcal{}L}_{\text{erosion}}$ is applied only to the centers of selected Gaussian pivots, not to all tetrahedral vertices, preventing collapse. This ensures regularization is effective without harming mesh integrity.
@@
To prevent such artifacts inside the mesh, we introduce a novel loss based on a feedback loop between the SDF values and the extracted mesh. This loss aims to enforce occluded points located inside the mesh to have negative SDF values. Specifically, after extracting the surface using the Gaussians and their SDF values as described in section 4.3, we use the mesh to build an occupancy label $o_{p}\in\{0,1\}$ for each Delaunay site $p$, indicating whether these points are inside or outside the visible portion of the surface mesh as described below. We finally enforce the Delaunay vertices labeled as inside to have negative SDF values, without enforcing any additional constraint for Delaunay vertices labeled as outside—as these points already receive supervision from the renderings.
@@
Figure 5. Qualitative comparison of surface reconstruction results on the Tanks and Temples dataset. MILo produces meshes of full scenes that better fit the target surfaces, with reduced artifacts such as cavities (three left images). Our approach effectively addresses the erosion problem that plagues previous methods (two right images), while producing lighter meshes. In addition, it is particularly effective at recovering peripheral objects (e.g., the bike rack or chandelier) that are often missed by post-hoc methods.
@@
implicit
NeRF (Mildenhall et al., 2020) 1.90 1.60 1.85 0.58 2.28 1.27 1.47 1.67 2.05 1.07 0.88 2.53 1.06 1.15 0.96 1.49 $\text{12h}$
VolSDF (Yariv et al., 2021) 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86 $\text{12h}$
NeuS (Wang et al., 2021) 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84 $\text{12h}$
Neuralangelo (Li et al., 2023) 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61 $\text{12h}$
explicit
3D GS (Kerbl et al., 2023) 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96 5.2m